Stock price = £30. In 2 months, two months the price will be either £33 or £27. The risk-free interest rate is 10% p.a on a continuous compounding basis.

What will be the value of a 2-month European put option with a strike price of £31? (5 marks)

Please provide a step by step explanation as I would like to fully understand and not just copy the answer. Thank you :)
Answer #1

Put option Payoff Cat =31) P £33 M So= £30 T-P £27 27 37 = 4 30 e 0.1 x 2112 33P+ 27(1-P) 6P = 30.5042- 27 P= 0.584 I-P= 0.41

Please do rate me and mention doubts in the comments section.